OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). Select one: a. -21 b. -3 c. 3 d. 21

OpenStudy (austinl):

Plug in -1 for all instances of x in g(x) Then take that value and plug it in for all instances of x in f(x).

OpenStudy (anonymous):

so -2*2-7*-2-9?

OpenStudy (austinl):

\(g(x)=x^2-7x-9\) \(g(-1)=(-1)^2-7(-1)-9=~?\)

OpenStudy (anonymous):

18?

OpenStudy (austinl):

Nope, you should get -1.

OpenStudy (anonymous):

how is that?

OpenStudy (austinl):

\((-1)^2=1\) \(-7(-1)=-(-7)=7\) \(1+7-9=8-9=-1\)

OpenStudy (anonymous):

ok then i take f(-1)=-1-2?

OpenStudy (austinl):

Yep!

OpenStudy (anonymous):

=3?

OpenStudy (austinl):

should be -3

OpenStudy (anonymous):

but then you multiply it by -1 and it makes it a positive right

OpenStudy (austinl):

x-2 x = -1 -1-2=-3

OpenStudy (anonymous):

but what about the f(-1)

OpenStudy (austinl):

f(-1) means that you just replace all instances of x with -1 in the function.

OpenStudy (anonymous):

oh ok

OpenStudy (austinl):

Good job!